# Resolvent-Based Optimisation for Invariant Flows

### v1

@@ 6 Conditioning of the Optimisation
Figure 15(a) shows the global residual traces for the perturbed field using $8$, $16$, $32$, and $64$ resolvent modes. All the residuals reduce at a high rate initially, indicating the power of the variational optimiser to quickly seek out the solution primarily by modifying the large-scale structures. This initially fast decrease, however, eventually gives way to a slower convergence rate, as is typically observed in the literature (Azimi et al., 2022; Ashtari & Schneider, 2023). The primary point to note from figure 15(a) is that both the degree of the eventual slowdown in convergence rate and the iteration count at which it occurs are strongly linked to the number of modes used for the Galerkin projection. The smaller the number of modes used the faster the overall convergence rate is, achieving the minimum residual in orders of magnitude fewer iterations than when more modes are retained. Specifically, the case of $M=64$ resolvent modes achieved its final residual after roughly $10^{5}$ iterations, whereas the $M=8$ case shows the same residual after only hundreds of iterations. The figure clearly demonstrates the truncation of the resolvent modes, derived about the laminar base flow, associated with the smallest singular values improves convergence. This provides confirmation that the condition number of the Hessian at the minimum is improved, even though the resolvent modes are derived relative to a different base flow. A final note is that an immediate result of truncating the resolvent modes is to reduce the global residual of the initial field. This is a result of the noise introduced in the perturbation of S3 being truncated in the projection, reducing the smaller scale noise present in the starting flow.
@@ 7 Conclusions
This work makes two primary contributions to the variational optimisation of exact coherent structures. First, we introduce a methodology for treating wall-bounded periodic flows by projecting the optimisation onto a basis of divergence-free resolvent modes that inherently enforce no-slip boundary conditions and incompressibility. This approach, which extends the methods of Farazmand (2016) and Azimi et al. (2022), directly embeds the required constraints, alleviating a key theoretical difficulty. This is in contrast to the approach taken in Ashtari & Schneider (2023) using the influence matrix method. In doing so, it unifies the optimisation framework with the resolvent-based modelling framework of McKeon & Sharma (2010), providing a direct method to “close the loop” in resolvent analysis (Barthel et al., 2021) by solving for the self-consistent, finite-amplitude velocity field that sustains the chosen forcing modes. Second, we establish a formal link between the conditioning of the optimisation problem and resolvent analysis via the Hessian operator. We demonstrate that truncating the resolvent basis not only creates a reduced-order model but also acts as an effective preconditioner thereby significantly accelerating convergence.
The projected optimisation methodology is applied to 2-dimensional, 3-component rotating plane Couette flow, with all analysis performed with a rotation number of $\mathit{Ro}=0.5$. Both equilibria and periodic solutions are sought in this work. The optimiser successfully identified equilibria that are not observed in DNS at $\mathit{Re}=50$ from various initial conditions. At $\mathit{Re}=400$, a stable periodic solution was obtained by initialising the optimiser with a perturbed DNS solution. In this case the residual decreased from $\mathcal{R}>10^{-3}$ to $\mathcal{R}<10^{-12}$ in roughly $40,000$ iterations, over half of which occurred in the first $5,000$. The slower convergence for the optimisation of the periodic solution is argued to be primarily a result of the increase in the degrees-of-freedom of the system and the possible weak linear stability of the solution.
@@
The projected variational optimiser presented in this work has two primary limitations. The most immediate is its high memory requirement for storing the high-dimensional state vector, which can become prohibitive for extending this method to fully three-dimensional turbulent flows. A promising path to mitigate this is leveraging parallel computing architectures, for which this framework is exceptionally well-suited since unlike traditional DNS the optimisation can be parallelised in the temporal dimension in addition to the spatial ones. A highly parallel implementation of the methodology would also permit an investigation into so-called “quasi-trajectories” from Burton et al. (2025) for turbulent flows. The second, more fundamental limitation is the efficacy of the resolvent basis being intrinsically linked to the chosen base flow. Determining a general principle for selecting the most effective base flow to yield useful, dynamically relevant modes requires extra investigation and represents a crucial direction for future work.
@@ A.1 Velocity Field Derivative
$\min_{\bm{u}\in\mathcal{P}_{\text{BC}},\,T}\quad\mathcal{R}\left[\bm{u},\,q\right]=\frac{1}{2}\norm{\bm{r}}^{2}_{\Omega_{t}}+\left\langle q,\,\divergence{\bm{u}}\right\rangle_{\Omega_{t}},$ (39)
@@
is the space of velocity fields that obey the boundary conditions, but aren’t necessarily incompressible. Also define the space of fields that obey the homogeneous boundary conditions
@@
where $\hat{\mathbf{n}}$ is an outwardly pointing normal vector on the surface $\partial\Omega$, $\differential{V} and d$$\differential{S}$ are infinitesimal volume and surface elements, respectively. Note that for any field $\bm{u}\in\mathcal{P}_{\text{HBC}}$ the surface integral vanishes, i.e.
@@
Using the fact that $\bm{v}\in\mathcal{P}_{\text{HBC}}$ and $\bm{u}\in\mathcal{P}_{\text{BC}}$, the surface integral vanish, leaving the following
@@
where the product rule and divergence theorem have been used twice to effectively integrate by parts the second-order derivatives in the Laplacian. The second of the boundary integrals vanish due to $\bm{v}\in\mathcal{P}_{T,\text{HMB}}$, however, the other boundary integral does not by default since we have not imposed any extra restrictions on the gradients of $\bm{v}$. To resolve this, it is necessary to impose the $\bm{r}\in\mathcal{P}_{T,\text{HBC}}$, enforcing that $\bm{r}$ vanishes at the walls, which means the first boundary integral also vanishes, giving
@@
Combining all of these results and substituting them into the right-hand side of (47) provides
@@
with the additional constraints $\bm{r}\in\mathcal{P}_{\text{HBC}}$ and $\divergence{\bm{r}}$. The boundary constraints on $\bm{r}$ at the boundaries is an instance of a natural boundary conditions, and are necessary to enforce. If they are not strictly enforced, then the residual gradient given in (61) is not guaranteed to be a descent direction. A similar interpretation applies to $\divergence{\bm{r}}=0$ except it applies throughout the whole domain rather than just the boundaries.
@@ A.2 Fundamental Frequency Derivative
The derivative of $\mathcal{R}$ with respect to the fundamental frequency $\omega$ is simpler to derive due to it not being functional and requiring methods from variational calculus. First, define a new scaled time $t^{\prime}=\omega t$ which then leads to a the following relationship between the time derivatives $\partialderivative*{t}=\omega\partialderivative*{t^{\prime}}$. Using this scaled time in the definition of the local residual in (5), and then substitute the resulting expression into the definition of the global residual
@@ Appendix C Resolvent for 2D3C RPCF
In this work we use resolvent analysis to generate an divergence-free and no-slip orthonormal basis set for the velocity and local residual fields. The exact form of the operators defined in section 3 for 2D3C RPCF is given here. First, the (nonlinear) Navier-Stokes operator is defined as
@@ Appendix D DNS Numerical Methods
The velocity vector field is conveniently decomposed into the laminar base state, and the velocity fluctuation denoted in this appendix as $\bm{u}(y,\,z,\,t)=(u\quad v\quad w)^{\top}$. A formulation using the streamwise component of the vorticity perturbation $\omega_{x}=\partial w/\partial y-\partial v/\partial z$, and the streamwise (fluctuation) velocity $u$, is used to numerically solve the problem (30). This formulation is used to eliminate complications arising from the pressure term. Because the domain is two-dimensional the stream function $\psi$, related to the transverse and wall-normal (fluctuation) velocity components by the relations $w=-\partial\psi/\partial y$ and $v={\partial\psi}/{\partial z}$, and to the vorticity by the Poisson-type equation $\bm{\Delta}\psi=-\omega$, is introduced. Introduction of the streamfunction ensure that the continuity equation is satisfied. With these definitions, the coupled system of nonlinear Partial Differential Equations (PDEs)

### v2

@@ 6 Conditioning of the Optimisation
Figure 14(a) shows the global residual traces for the perturbed field using $8$, $16$, $32$, and $64$ resolvent modes. All the residuals reduce at a high rate initially, indicating the power of the variational optimiser to quickly seek out the solution primarily by modifying the large-scale structures. This initially fast decrease, however, eventually gives way to a slower convergence rate, as is typically observed in the literature (Azimi et al., 2022; Ashtari & Schneider, 2023). The primary point to note from figure 14(a) is that both the degree of the eventual slowdown in convergence rate and the iteration count at which it occurs are strongly linked to the number of modes used for the Galerkin projection. The smaller the number of modes used the faster the overall convergence rate is, achieving the minimum residual in orders of magnitude fewer iterations than when more modes are retained. Specifically, the case of $M=64$ resolvent modes achieved its final residual after roughly $10^{5}$ iterations, whereas the $M=8$ case shows the same residual after only hundreds of iterations. The figure clearly demonstrates the truncation of the resolvent modes, derived about the laminar base flow, associated with the smallest singular values improves convergence. This provides confirmation that the condition number of the Hessian at the minimum is improved, even though the resolvent modes are derived relative to a different base flow. A final note is that an immediate result of truncating the resolvent modes is to reduce the global residual of the initial field, marked on figure 14 at the start of each optimisation. This is a result of the noise introduced in the perturbation of S3 being truncated in the projection, reducing the smaller scale noise present in the starting flow.
@@ 7 Conclusions
This work makes two primary contributions to the variational optimisation of exact coherent structures. First, we introduce a methodology for treating wall-bounded time periodic flows by projecting the optimisation onto a basis of divergence-free resolvent modes that inherently enforce no-slip boundary conditions and incompressibility. This approach, which extends the methods of Farazmand (2016) and Azimi et al. (2022), directly embeds the required constraints, alleviating a key theoretical difficulty. This is in contrast to the approach taken in Ashtari & Schneider (2023) using the influence matrix method. In doing so, it unifies the optimisation framework of these earlier works for finding ECSs with the resolvent-based modelling framework of McKeon & Sharma (2010), providing a direct method to “close the loop” in resolvent analysis (Barthel et al., 2021) by solving for the self-consistent, finite-amplitude velocity field that sustains the chosen forcing modes. Second, we establish a formal link between the conditioning of the optimisation problem and resolvent analysis via the Hessian operator. We demonstrate that truncating the resolvent basis not only acts as an effective preconditioner for the optimiser, but also benefits from the resulting reduced-order model reducing the dimensionality of the problem while still leading to physically meaningful results.
The projected optimisation methodology is applied to 2-dimensional, 3-component rotating plane Couette flow, with all analysis performed with a rotation number of $\mathit{Ro}=0.5$. Both equilibria and periodic solutions are sought in this work. To ensure only minimal a priori knowledge about the flow is required the laminar profile is used to generate the required resolvent modes instead of the turbulent mean as is more standard for typical resolvent analysis applications. The implications of this choice and the effect it has on the modes has not been investigated here and warrants further study. The optimiser successfully identified equilibria that are not observed in DNS at $\mathit{Re}=50$ from various initial conditions. At $\mathit{Re}=400$, a stable periodic solution was obtained by initialising the optimiser with a perturbed DNS solution. In this case the residual decreased from $\mathcal{R}>10^{-3}$ to $\mathcal{R}<10^{-12}$ in roughly $40,000$ iterations, over half of which occurred in the first $5,000$. The slower convergence for the optimisation of the periodic solution is argued to be primarily a result of the increase in the degrees-of-freedom of the system and the possible weak linear stability of the solution.
@@
The projected variational optimiser presented in this work has two primary limitations. The most immediate is its high memory requirement for storing the high-dimensional state vector, which can become prohibitive for extending this method to fully three-dimensional turbulent flows. A promising path to mitigate this is leveraging parallel computing architectures, for which this framework is exceptionally well-suited since, unlike traditional direct numerical simulation of turbulence, the optimisation can be parallelised in the temporal dimension in addition to the spatial ones. A highly parallel implementation of the methodology would also permit an investigation into so-called “quasi-trajectories” from Burton et al. (2025) for turbulent flows. These quasi-trajectories would constitute very low-order representations of the flow, similar in spirit to those explored by Li & Lasagna (2025). However, whereas Li & Lasagna (2025) considered cases where the final objective was driven to very small values, sometimes near machine precision, the quasi-trajectories discussed by Burton et al. (2025) would not aim to satisfy the low-order projection of the governing equations to such an extent. The second, more fundamental limitation is the efficacy of the resolvent basis being intrinsically linked to the chosen base flow. Determining a general principle for selecting the most effective base flow to yield useful, dynamically relevant modes requires extra investigation and represents an important direction for future work.
{bmhead}
[Supplementary movies.] Supplementary movies are available at
{bmhead}
[Funding.] This work was supported using the Antony Wright Scholarship provided by the Department of Aeronautics and Astronautics at the University of Southampton.
{bmhead}
[Competing interests.] The authors declare none.
{bmhead}
[Author ORCIDs.
] Thomas Burton https://orcid.org/0000-0001-7998-2278;
Sean Symon https://orcid.org/0000-0001-9085-0778;
Davide Lasagna https://orcid.org/0000-0002-6501-6041;
@@ A.1 Velocity Field Derivative
$\min_{\bm{u}\in\mathcal{P}_{T,\text{BC}},\,T}\quad\mathcal{R}\left[\bm{u},\,q\right]=\frac{1}{2}\norm{\bm{r}}^{2}_{\Omega_{t}}+\left\langle q,\,\divergence{\bm{u}}\right\rangle_{\Omega_{t}},$ (39)
@@
is the space of velocity fields that obey the boundary conditions, but are not necessarily incompressible. Also define the space of fields that obey the homogeneous boundary conditions
@@
where $\hat{\mathbf{n}}$ is an outwardly pointing normal vector on the surface $\partial\Omega$, $\differential{V} and d$$\differential{S}$ are infinitesimal volume and surface elements, respectively. Note that for any field $\bm{u}\in\mathcal{P}_{T,\text{HBC}}$ the surface integral vanishes, i.e.
@@
Using the fact that $\bm{v}\in\mathcal{P}_{T,\text{HBC}}$ and $\bm{u}\in\mathcal{P}_{T,\text{BC}}$, the surface integral vanish, leaving the following
@@
where the product rule and divergence theorem have been used twice to effectively integrate by parts the second-order derivatives in the Laplacian. The second of the boundary integrals vanish due to $\bm{v}\in\mathcal{P}_{T,\text{HMB}}$, however, the other boundary integral does not by default since we have not imposed any extra restrictions on the gradients of $\bm{v}$. To resolve this, it is necessary to impose $\bm{r}\in\mathcal{P}_{T,\text{HBC}}$, enforcing that $\bm{r}$ vanishes at the walls, which means the first boundary integral also vanishes, giving
@@
Combining all of these results and substituting them into the right-hand side of (47) results in
@@
with the additional constraints $\bm{r}\in\mathcal{P}_{T,\text{HBC}}$ and $\divergence{\bm{r}}=0$. The boundary constraints on $\bm{r}$ at the boundaries is an instance of a natural boundary conditions, and are necessary to enforce. If they are not strictly enforced, then the residual gradient given in (61) is not guaranteed to be a descent direction. A similar interpretation applies to $\divergence{\bm{r}}=0$ except it applies throughout the whole domain rather than just the boundaries.
@@ A.2 Fundamental Frequency Derivative
The derivative of $\mathcal{R}$ with respect to the fundamental frequency $\omega$ is simpler to derive since $\mathcal{R}$ is an ordinary function of $\omega$ rather than a functional such as with $\bm{u}$. First, define a new scaled time $t^{\prime}=\omega t$ which then leads to a the following relationship between the time derivatives $\partialderivative*{t}=\omega\partialderivative*{t^{\prime}}$. Using this scaled time in the definition of the local residual in (5), and then substituting the resulting expression into the definition of the global residual gives
@@ Appendix C Resolvent for 2D3C RPCF
In this work we use resolvent analysis to generate a divergence-free and no-slip orthonormal basis set for the velocity and local residual fields. The exact form of the operators defined in section 3 for 2D3C RPCF is given here. First, the (nonlinear) Navier-Stokes operator is defined as
@@ Appendix D DNS Numerical Methods
The velocity vector field is conveniently decomposed into the laminar base state, and the velocity fluctuation denoted in this appendix as $\bm{u}(y,\,z,\,t)=(u\quad v\quad w)^{\top}$. A formulation using the streamwise component of the vorticity perturbation $\omega_{x}=\partial w/\partial y-\partial v/\partial z$, and the streamwise (fluctuation) velocity $u$, is used to numerically solve the problem (30). This formulation eliminates complications arising from the pressure term. Because the domain is two-dimensional the stream function $\psi$, related to the transverse and wall-normal (fluctuation) velocity components by the relations $w=-\partial\psi/\partial y$ and $v={\partial\psi}/{\partial z}$, and to the vorticity by the Poisson-type equation $\bm{\Delta}\psi=-\omega$, is introduced. Introduction of the streamfunction ensures that the continuity equation is satisfied. With these definitions, the coupled system of nonlinear Partial Differential Equations (PDEs)